Can someone help me on this problem?
I need to solve for x and check my result:
3/7x = 15
Thanks,
Sher

Answer by checkley71(8403)

IF THIS IS (3/7)X=15 THEN 3X=105 OR X=35
PROOF 3/7*35=15 OR 3*5=35 OR 35=35
IF THIS IS 3/(7X)=15 THEN 105X=3 OR X=3/105 OR X=1/35
3/(7(1/35))=15 OR 3/(1/5)=15 OR 3*5=15 OR 15=15